OpenStudy (anonymous):

What is 16% of 25? A. 64 B. 16 C. 6 D. 4

OpenStudy (anonymous):

Looks like same question as knaim"s

OpenStudy (anonymous):

what???

OpenStudy (anonymous):

whatever.. There's a quick and easy way to do this: let's pretend the 25 is 100%, and the number we're looking for is x Therefore: x/25 = 16/100 the 16 corresponds to the number you're looking for, and the 100 is the 25. Now we simply cross multiply: 100x/25 = 16 4x = 16 x=4

OpenStudy (shamim):

25*16%=25*16/100

OpenStudy (anonymous):

The other way to do this is to divide 25 by 100 = 0.25 which means 0.25 is 1% so 0.25 x 16 = 4 so that means that 4 is 16 % of 25

OpenStudy (anonymous):

so shamim's equation is correct

OpenStudy (anonymous):

25/100 =.25 so 16% from 25 equal 0.25 *16=4

OpenStudy (anonymous):

hanishala is also right

OpenStudy (anonymous):

thanks